OpenStudy (anonymous):

lim 3/e^x + 7 x->INF How do I solve this?

OpenStudy (anonymous):

3/e^x tends to 0 as x tends to infinity So your lim=0+7=7

OpenStudy (zarkon):

Andras' solution is correct the way you typed the problem, but did you intend for the problem to be \[\lim_{x\to\infty}\frac{3}{e^x+7}\] ?

OpenStudy (anonymous):

yea that's the prob i meant

OpenStudy (zarkon):

then the answer is 0 since \[e^x+7\to\infty\text{ as }x\to\infty\]

OpenStudy (anonymous):

thank ya